OpenStudy (anonymous):

if a guy had $100 and he wanted to buy a pair of nice nikes and they were $60.99 but he also saw these really nice converese for $39.00 he bought both of those now how much money would he have left if he went to another store and bought a pair of sock for $3.00 how much monery does he have left ?

OpenStudy (anonymous):

sorry for the misspeled words gosh :)

OpenStudy (anonymous):

100-60.99=39.01 39.01-39=0.01 0.01-3=-2.99 He can only do this if he gets $2.99 more than the $100 he already has

OpenStudy (anonymous):

okay thank you guys so much for answering :)

OpenStudy (anonymous):

He would have -2.99

OpenStudy (anonymous):

:) ty